2 Consider the ANOVA table that follows. Analysis of Variance Source Regression Residual Error47 2,139.35 Total DF MS 12.5 points 6 3,620.71 603.45 13.26 45.52 53 5,760.06 Skipped eBooka-1. Determine the standard error of estimate. (Round your answer to 2 decimal places.) References Standard error of estimate a-2. About 95% of the residuals will be between what two values? (Round your answers to 2 decimal places.) 95% of the residuals will be between an b-1. Determine the coefficient of multiple determination. (Round your answer to 3 decimal places.)2 b-1. Determine the coefficient of multiple determination. (Round your answer to 3 decimal places.) 2.5 points Coefficient of multiple determination Skipped eBook b-2. Determine the percentage variation for the independent variables. (Round your answer to 1 decimal place.) References he independent vari % of the variation ain c. Determine the coefficient of multiple determination, adjusted for the degrees of freedom. (Round your answer to 3 decimal places.) Coefficient of multiple determination
Answer #1

a-1)

std error of estimate =sqrt(MSE)=sqrt(45.52)=6.75

a-2)

95% of residuals will be between -13.5 and 13.5

b-1)

coefficient of deternination =R2 =SSR/SST=0.629

b-2)

% of variation =r2*100 =62.9%

c)

R2adj= 1-(n-1)*SSE/(SST(n-k-1)) = 0.581